Superior Division of the Monroe Company has to invest in a new project. The project an incremental operating income ol $73.350 on average invested assets of $900,000. Superior Division currenty has operating income of $425,000 on average invested $4,325,000,Monroe Company has a 7% hurdle ratefor new projects. (25 points) assets of a. What is Superior Divisions ROI before making an investment in the project? 42 5oco olz?25,000. b. What is Superior Divisions residual income before making an investment in the project? 925/000- ( 325000 x %) oco- 302, 75 122,25o R c. What is Superior Divisions ROl after making the investment in the project? 325090 522pa d. What is Superior Divisions residual income after making the investment in the project? 98,350 365,A5 - 132,6o0
Answer #1
a
ROI = 4325000/425000= 10.18%
b
Residual income = 425000-(4325000*7%)= $122250
c
ROI = (425000+73350)/(4325000+900000)= 9.54%
d
Residual income = (425000+73350)-(4325000+900000)*7%= $132600